Please solve the equation using the zero-factor property. Thank you
5x^2+12=23x

Answer by checkley75(3666)

5X^2-23X+12=0
(5X-3)(X-4)=0
5X-3=0
5X=3
X=3/5 ANSWER.
X-4=0
X=4 ANSWER.